OpenStudy (anonymous):

x+5(x+3)=27 whats x?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

x +5x + 15 = 27 6x +15 = 27 -15 -15 6x = 12 /6 /6 x=2